The following sample shows the shirt colors of a 
random sample of students in school. Find the 
experimental probability a student chosen at random 
is wearing a white shirt. 
blue 8
red 7
white 12
other 3


Answer by AnlytcPhil(1806)


The following sample shows the shirt colors
of a random sample of students in school. 
Find the experimental probability a student
chosen at random is wearing a white shirt. 
blue 8
red 7
white 12
other 3

Total in sample = 8 + 7 + 12 + 3 = 30

P(blue) = 8 out of 30 = 8/30 = 4/15
P(red) = 7 out of 30 = 7/30
P(white) = 12 out of 30 = 12/30 = 2/5
P(other) = 3 out of 30 = 3/30 = 1/10

Answer:  P(white) = 2/5 

Edwin McCravy
AnlytcPhil@aol.com